THE CAVITY RING–DOWN ABSORPTION SPECTRUM OF THE S_{0} → T_{1} AND S_{0} → S_{1} TRANSITION OF JET–COOLED AROMATIC THIONE COMPOUNDS

\underline{A.A. RUTH}^{a}, Institut für Atomare und Analytische Physik, TU–Berlin, Hardenbergstr. 36, 10623 Berlin, Germany; T. FERNHOLZ, Physikalisch Chemisches Institut, Universität Heidelberg, Im Neuenheimer Feld 253, 69120 Heidelberg, Germany.

The cavity ring–down method has been applied to the measurement of the absorption of 4–H–1–benzopyrane–4–thione (BPT)^{b} and 4–H–pyrane–4–thione (PT) in a supersonic jet in the wavelength region of the weak S_{0} → T_{1} and S_{0} → S_{1} transition. In the case of jet–cooled BPT the absorption energies of vibronic states between 15800 and 16650 cm^{−1} corroborate previous assignments in the S_{0} → T_{1} phosphorescence excitation and resonance–enhanced multi–photon excitation spectra. The symmetry forbidden 0,0 transition, S_{0,0} → S_{1,0} at 16522 cm^{−1}, was found to be 7.5 times weaker than the absorption transition to the triplet origin, S_{0,0} → T_{1z,0}. Relative phosphorescence quantum yields were estimated. For the more volatile PT the CRD absorption spectrum of jet–cooled PT was compared to PT in a static cell. Through the hot band analysis ground state vibrations were assigned in comparison to ab initio force field calculations^{c}

aSupport by the DFG (RU672/1-1/1-2) is gratefully acknowledged

bA.A. Ruth, T. Fernholz, R.P. Brint and M.W.D. Mansfield; Chem. Phys. Lett. (1998) in press.

cÁ. Somogyi, G. Jalsovszky, C. Fülöp, J. Stark and J.E. Boggs; Spectrochimica Acta \underline{45a} (1989) 679.